# Sample-Efficient Reinforcement Learning from Human Feedback via Information-Directed Sampling

### v1

@@ A-A Proof of Theorem 1
Therefore, by Lemma 4, we have
$\mathbb{E}_{t}\left[V_{1,\pi_{\mathcal{E}}^{*}}^{\tilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})-V_{1,\pi_{\text{TS}}^{t}}^{\tilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})\right]$ $=\mathbb{E}_{t}\left[V_{1,\pi_{\mathcal{E}}^{*}}^{\tilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})-V_{1,\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}(s_{1}^{t})\right]$
$=\sum_{h=1}^{H}\mathbb{E}_{t}\mathbb{E}_{\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}\left[\Delta_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(s,a)\right]$
$=\sum_{h=1}^{H}\mathbb{E}_{t}\left[\int_{\mathcal{S}\times\mathcal{A}}d_{h,\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}(s,a)\Delta_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(s,a)\mathrm{d}\mu_{\mathcal{S}\times\mathcal{A}}\right],$ (33)
where the notation of $d_{h,\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}(s,a)$ and $\Delta_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(s,a)$ are introduced to simplify the formula. Following [7], we define
$\mathcal{I}^{t}\triangleq\sum_{h=1}^{H}\mathbb{E}_{t}\mathbb{E}_{\pi_{\mathrm{TS}}^{t}}^{\bar{\mathcal{E}}_{t}}\left[\frac{\Delta_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(s,a)^{2}}{\alpha_{\mathcal{E}}^{2}}\right],\quad\mathcal{T}^{t}\triangleq\sum_{h=1}^{H}\int_{\mathbb{E}_{t}[d_{h,\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}(s,a)]\neq 0}\frac{\mathbb{E}_{t}\left[\alpha_{\mathcal{E}}^{2}\cdot d_{h,\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}(s,a)^{2}\right]}{\mathbb{E}_{t}\left[d_{h,\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}(s,a)\right]}\mathrm{d}\mu_{\mathcal{S}\times\mathcal{A}}.$ (34)
@@
$\sum_{h=1}^{H}\mathbb{E}_{t}\left[\int_{\mathcal{S}\times\mathcal{A}}d_{h,\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}(s,a)\Delta_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(s,a)\mathrm{d}\mu_{\mathcal{S}\times\mathcal{A}}\right]$
$=\sum_{h=1}^{H}\mathbb{E}_{t}\left[\int_{\mathbb{E}_{t}[d_{h,\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}(s,a)]\neq 0}d_{h,\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}(s,a)\Delta_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(s,a)\mathrm{d}\mu_{\mathcal{S}\times\mathcal{A}}\right]$
$\leq\left(\sum_{h=1}^{H}\mathbb{E}_{t}\int_{\mathbb{E}_{t}[d_{h,\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}(s,a)]\neq 0}\frac{\alpha_{\mathcal{E}}^{2}\cdot d_{h,\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}(s,a)^{2}}{\mathbb{E}_{t}\left[d_{h,\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}(s,a)\right]}\right)^{1/2}\left(\sum_{h=1}^{H}\mathbb{E}_{t}\int_{\mathcal{S}\times\mathcal{A}}\mathbb{E}_{t}\left[d_{h,\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}(s,a)\right]\cdot\frac{\Delta_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(s,a)^{2}}{\alpha_{\mathcal{E}}^{2}}\right)^{1/2}$
$=\sqrt{\mathcal{I}^{t}\cdot\mathcal{T}^{t}},$ (35)
where the first equality is due to the fact that $\Delta_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(s,a)$ is bounded ($\leq 2H$), and the second inequality is simply the Cauchy-Schwarz inequality with $\sum_{h}\mathbb{E}_{t}\int_{\mathcal{S}\times\mathcal{A}}$ as an “integrated” integral over the space $[H]\times\Theta\times\mathcal{S}\times\mathcal{A}$. Let us briefly discuss why the third equality holds. For the term $\mathcal{T}^{t}$, the derivation is straightforward, since $\mathbb{E}_{t}[X/\mathbb{E}_{t}[Y]]=\mathbb{E}_{t}[X]/\mathbb{E}_{t}[Y]$. For the term $\mathcal{I}^{t}$, first recall that $\mathbb{E}_{t}\left[d_{h,\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}(s,a)\right]=\mathbb{E}_{t}\left[d_{h,\pi_{\mathrm{TS}}^{t}}^{\bar{\mathcal{E}}_{t}}(s,a)\right]$ due to the property of TS. Then, we can use the independence between $d_{h,\pi_{\mathrm{TS}}^{t}}^{\bar{\mathcal{E}}_{t}}(s,a)$ and $\Delta_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(s,a)$ given $\mathcal{D}_{t}$ to “extract” the expectation ($\mathbb{E}[XY]=\mathbb{E}[X]\cdot\mathbb{E}[Y]$ for $X,Y$ mutually independent):
$\sum_{h=1}^{H}\mathbb{E}_{t}\int_{\mathcal{S}\times\mathcal{A}}\mathbb{E}_{t}\left[d_{h,\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}(s,a)\right]\cdot\frac{\Delta_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(s,a)^{2}}{\alpha_{\mathcal{E}}^{2}}=\sum_{h=1}^{H}\mathbb{E}_{t}\int_{\mathcal{S}\times\mathcal{A}}d_{h,\pi_{\mathrm{TS}}^{t}}^{\bar{\mathcal{E}}_{t}}(s,a)\cdot\frac{\Delta_{h}^{\tilde{\mathcal{E}}_{t}^{*}}(s,a)^{2}}{\alpha_{\mathcal{E}}^{2}}=\mathcal{I}^{t}.$ (36)
To summarize, by Eqn. (35) we have the following bound for $\Gamma_{t}^{\pi_{\mathrm{TS}}^{t}}$:
$\Gamma_{t}^{\pi_{\mathrm{TS}}^{t}}\leq\frac{\mathcal{I}^{t}\cdot\mathcal{T}^{t}}{\mathbb{I}_{t}^{\pi_{\mathrm{TS}}^{t}}\left(\tilde{\mathcal{E}}_{t}^{*};(\mathcal{H}_{t},\mathcal{R}_{t,H})\right)}.$ (37)
@@
$\mathcal{I}^{t}$ $=\sum_{h=1}^{H}\mathbb{E}_{t}\mathbb{E}_{\pi_{\text{TS}}^{t}}^{\bar{\mathcal{E}}_{t}}\bigg{(}\mathbb{E}_{(s^{\prime},r^{\prime})\sim(P_{h}^{\tilde{\mathcal{E}}_{t}^{*}}\otimes R_{h}^{\tilde{\mathcal{E}}_{t}^{*}})(\cdot|s_{h},a_{h})}\bigg{[}\frac{r^{\prime}+V_{h+1,\pi_{\mathcal{E}}^{*}}^{\tilde{\mathcal{E}}_{t}^{*}}(s^{\prime})}{\alpha_{\mathcal{E}}}\bigg{]}-\mathbb{E}_{(s^{\prime},r^{\prime})\sim(P_{h}^{\bar{\mathcal{E}}_{t}}\otimes R_{h}^{\bar{\mathcal{E}}_{t}})(\cdot|s_{h},a_{h})}\bigg{[}\frac{r^{\prime}+V_{h+1,\pi_{\mathcal{E}}^{*}}^{\tilde{\mathcal{E}}_{t}^{*}}(s^{\prime})}{\alpha_{\mathcal{E}}}\bigg{]}\bigg{)}^{2}$
$=\sum_{h=1}^{H}\mathbb{E}_{t}\mathbb{E}_{\pi_{\text{TS}}^{t}}^{\bar{\mathcal{E}}_{t}}\bigg{(}\mathbb{E}_{(s^{\prime},r^{\prime})\sim(P_{h}^{\tilde{\mathcal{E}}_{t}^{*}}\otimes R_{h}^{\tilde{\mathcal{E}}_{t}^{*}})(\cdot|s_{h},a_{h})}\bigg{[}\frac{r^{\prime}-r_{h}^{\inf}(s_{h},a_{h})+V_{h+1,\pi_{\mathcal{E}}^{*}}^{\tilde{\mathcal{E}}_{t}^{*}}(s^{\prime})-\inf_{s}V_{h+1,\pi_{\mathcal{E}}^{*}}^{\tilde{\mathcal{E}}_{t}^{*}}(s^{\prime})}{\alpha_{\mathcal{E}}}\bigg{]}$ (38)
$\qquad-\mathbb{E}_{(s^{\prime},r^{\prime})\sim(P_{h}^{\bar{\mathcal{E}}_{t}}\otimes R_{h}^{\bar{\mathcal{E}}_{t}})(\cdot|s_{h},a_{h})}\bigg{[}\frac{r^{\prime}-r_{h}^{\inf}(s_{h},a_{h})+V_{h+1,\pi_{\mathcal{E}}^{*}}^{\tilde{\mathcal{E}}_{t}^{*}}(s^{\prime})-\inf_{s}V_{h+1,\pi_{\mathcal{E}}^{*}}^{\tilde{\mathcal{E}}_{t}^{*}}(s^{\prime})}{\alpha_{\mathcal{E}}}\bigg{]}\bigg{)}^{2}$
$\leq\frac{1}{2}\sum_{h=1}^{H}\mathbb{E}_{t}\mathbb{E}_{\pi_{\text{TS}}^{t}}^{\bar{\mathcal{E}}_{t}}\left[D_{\text{KL}}\left(\left(P_{h}^{\tilde{\mathcal{E}}_{t}^{*}}\otimes R_{h}^{\tilde{\mathcal{E}}_{t}^{*}}\right)(\cdot|s_{h},a_{h})\middle\|\left(P_{h}^{\bar{\mathcal{E}}_{t}}\otimes R_{h}^{\bar{\mathcal{E}}_{t}}\right)(\cdot|s_{h},a_{h})\right)\right]$
$\leq\frac{1}{2}\mathbb{I}_{t}^{\pi_{\text{TS}}^{t}}\left(\tilde{\mathcal{E}}_{t}^{*};(\mathcal{H}_{t},\mathcal{R}_{t,H})\right),$ (39)
where the two inequalities are due to Pinsker’s inequality and Lemma 5 respectively. Plugging into Eqn. (37), we derive that $\Gamma_{t}^{\pi_{\text{TS}}^{t}}\leq\frac{1}{2}\mathcal{T}^{t}$, and thus
$BR_{T}(\pi_{\mathrm{IDS}})\leq\frac{1}{4\lambda}\mathbb{E}\left[\sum_{t=1}^{T}\mathcal{T}^{t}\right]+\frac{\lambda}{2}\log(K(\epsilon))+T\epsilon.$ (40)
Step 4 (Bound $\mathbb{E}[\mathcal{T}^{t}]$). The analysis tools used in this step is the Doob’s consistency theorem, with more details discussed in Appendix C-B. Define the true environment as $\mathcal{E}_{0}$. For brevity of notations, we define
@@
$\mathcal{T}^{t}=\sum_{h=1}^{H}\int_{(s,a)\in\mathcal{B}_{h,t}}\frac{\mathbb{E}_{t}\left[\alpha_{\mathcal{E}}^{2}\cdot d_{h,\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}(s,a)^{2}\right]}{\mathbb{E}_{t}\left[d_{h,\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}(s,a)\right]}\mathrm{d}\mu_{\mathcal{S}\times\mathcal{A}}.$ (41)
We now introduce another variable $\mathcal{E}^{\prime}$ to apply Lemma 6. Let $\mathcal{E}^{\prime}\sim\mathbb{P}(\cdot|\mathcal{D}_{t})$ be a random variable independent of $\mathcal{E}$. By definition of $d_{h,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}}$, we have $\mathbb{E}_{t}\left[d_{h,\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}(s,a)\right]=\mathbb{E}_{\mathcal{E},\mathcal{E}^{\prime}\sim\mathbb{P}_{t}(\cdot)}\left[d_{h,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}^{\prime}}(s,a)\right]$, and
@@
$\leq\mathbb{E}_{\mathcal{E},\mathcal{E}^{\prime}\sim\mathbb{P}_{t}(\cdot)}\left[\alpha_{\mathcal{E}}^{2}\cdot d_{h,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}^{\prime}}(s,a)^{2}\right],$ (42)
@@
$\mathcal{T}^{t}\leq\sum_{h=1}^{H}\int_{\mathcal{S}\times\mathcal{A}}\frac{\mathbb{E}_{t}\left[\alpha_{\mathcal{E}}^{2}\cdot d_{h,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}^{\prime}}(s,a)^{2}\right]}{\mathbb{E}_{t}\left[d_{h,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}^{\prime}}(s,a)\right]}\chi_{\mathcal{B}_{h,t}}(s,a)\mathrm{d}\mu_{\mathcal{S}\times\mathcal{A}}=\sum_{h=1}^{H}\int_{\mathcal{S}\times\mathcal{A}}g_{t}(s,a,h,\mathcal{D}_{t})\mathrm{d}\mu_{\mathcal{S}\times\mathcal{A}},$ (43)
@@
This implies $g_{t}(s,a,h,\mathcal{D}_{t})\leq M_{d}H^{2}$ and $\mathcal{T}^{t}\leq M_{d}H^{3}$. By Lemma 6, we have the following convergence results: for any $(s,a)\in\mathcal{S}\times\mathcal{A}$,
$\lim_{t\to\infty}\mathbb{E}_{t}\left[\alpha_{\mathcal{E}}^{2}d_{h,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}^{\prime}}(s,a)^{2}\right]=\alpha_{\mathcal{E}_{0}}^{2}d_{h,\pi_{\mathcal{E}_{0}}^{*}}^{\mathcal{E}_{0}}(s,a)^{2},$ (44)
$\lim_{t\to\infty}\mathbb{E}_{t}\left[d_{h,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}^{\prime}}(s,a)\right]=d_{h,\pi_{\mathcal{E}_{0}}^{*}}^{\mathcal{E}_{0}}(s,a),$ (45)
@@
$\lim_{t\to\infty}\chi_{\mathcal{B}_{h,t}}(x)=\chi_{\mathcal{B}_{h}}(x),\qquad\text{where }\mathcal{B}_{h}\triangleq\left\{(s,a)\in\mathcal{S}\times\mathcal{A}\middle|\mathbb{E}\left[d_{h,\pi_{\mathcal{E}_{0}}^{*}}^{\mathcal{E}_{0}}(s,a)\right]\neq 0\right\}.$ (46)
From Eqn. (46), we can also derive that $\lim_{t\to\infty}\chi_{\mathcal{B}_{h,t}\bigcap\mathcal{B}_{h}}(x)=1$, and $\lim_{t\to\infty}\chi_{\mathcal{B}_{h,t}\setminus\mathcal{B}_{h}}(x)=0$. By Lebesgue dominated convergence theorem, we have

### v3

@@ Appendix A Proofs of Theorem and Proposition
Therefore, by Lemma C.1, we have
$\mathbb{E}_{t}\left[V_{1,\pi_{\mathcal{E}}^{*}}^{\widetilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})-V_{1,\pi_{\text{TS}}^{t}}^{\widetilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})\right]$ $=\mathbb{E}_{t}\left[V_{1,\pi_{\mathcal{E}}^{*}}^{\widetilde{\mathcal{E}}_{t}^{*}}(s_{1}^{t})-V_{1,\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}(s_{1}^{t})\right]$
$=\sum_{h=1}^{H}\mathbb{E}_{t}\mathbb{E}_{\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}\left[\Delta_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(s,a)\right]$
$=\sum_{h=1}^{H}\mathbb{E}_{t}\left[\int_{\mathcal{S}\times\mathcal{A}}d_{h,\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}(s,a)\Delta_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(s,a)\mathrm{d}\mu_{\mathcal{S}\times\mathcal{A}}\right],$ (A.9)
where the notation of $d_{h,\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}(s,a)$ and $\Delta_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(s,a)$ are introduced to simplify the formula. Following Moradipari et al. (2023), we define
$\mathcal{I}^{t}\triangleq\sum_{h=1}^{H}\mathbb{E}_{t}\mathbb{E}_{\pi_{\mathrm{TS}}^{t}}^{\bar{\mathcal{E}}_{t}}\left[\frac{\Delta_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(s,a)^{2}}{\alpha_{\mathcal{E}}^{2}}\right],\quad\mathcal{T}^{t}\triangleq\sum_{h=1}^{H}\int_{\mathbb{E}_{t}[d_{h,\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}(s,a)]\neq 0}\frac{\mathbb{E}_{t}\left[\alpha_{\mathcal{E}}^{2}\cdot d_{h,\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}(s,a)^{2}\right]}{\mathbb{E}_{t}\left[d_{h,\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}(s,a)\right]}\mathrm{d}\mu_{\mathcal{S}\times\mathcal{A}}.$ (A.10)
@@
$\sum_{h=1}^{H}\mathbb{E}_{t}\left[\int_{\mathcal{S}\times\mathcal{A}}d_{h,\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}(s,a)\Delta_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(s,a)\mathrm{d}\mu_{\mathcal{S}\times\mathcal{A}}\right]$
$=\sum_{h=1}^{H}\mathbb{E}_{t}\left[\int_{\mathbb{E}_{t}[d_{h,\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}(s,a)]\neq 0}d_{h,\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}(s,a)\Delta_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(s,a)\mathrm{d}\mu_{\mathcal{S}\times\mathcal{A}}\right]$
$\leq\left(\sum_{h=1}^{H}\mathbb{E}_{t}\int_{\mathbb{E}_{t}[d_{h,\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}(s,a)]\neq 0}\frac{\alpha_{\mathcal{E}}^{2}\cdot d_{h,\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}(s,a)^{2}}{\mathbb{E}_{t}\left[d_{h,\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}(s,a)\right]}\right)^{1/2}\left(\sum_{h=1}^{H}\mathbb{E}_{t}\int_{\mathcal{S}\times\mathcal{A}}\mathbb{E}_{t}\left[d_{h,\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}(s,a)\right]\cdot\frac{\Delta_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(s,a)^{2}}{\alpha_{\mathcal{E}}^{2}}\right)^{1/2}$
$=\sqrt{\mathcal{I}^{t}\cdot\mathcal{T}^{t}},$ (A.11)
where the first equality is due to the fact that $\Delta_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(s,a)$ is bounded ($\leq 2H$), and the second inequality is simply the Cauchy-Schwarz inequality with $\sum_{h}\mathbb{E}_{t}\int_{\mathcal{S}\times\mathcal{A}}$ as an “integrated” integral over the space $H]\times\Theta\times\mathcal{S}\times\mathcal{A}$. Let us briefly discuss why the third equality holds. For the term $\mathcal{T}^{t}$, the derivation is straightforward, since $\mathbb{E}_{t}[X/\mathbb{E}_{t}[Y]]=\mathbb{E}_{t}[X]/\mathbb{E}_{t}[Y]$. For the term $\mathcal{I}^{t}$, first recall that $\mathbb{E}_{t}\left[d_{h,\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}(s,a)\right]=\mathbb{E}_{t}\left[d_{h,\pi_{\mathrm{TS}}^{t}}^{\bar{\mathcal{E}}_{t}}(s,a)\right]$ due to the property of TS. Then, we can use the independence between $d_{h,\pi_{\mathrm{TS}}^{t}}^{\bar{\mathcal{E}}_{t}}(s,a)$ and $\Delta_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(s,a)$ given $\mathcal{D}_{t}$ to “extract” the expectation ($\mathbb{E}[XY]=\mathbb{E}[X]\cdot\mathbb{E}[Y]$ for $X,Y$ mutually independent):
$\sum_{h=1}^{H}\mathbb{E}_{t}\int_{\mathcal{S}\times\mathcal{A}}\mathbb{E}_{t}\left[d_{h,\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}(s,a)\right]\cdot\frac{\Delta_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(s,a)^{2}}{\alpha_{\mathcal{E}}^{2}}=\sum_{h=1}^{H}\mathbb{E}_{t}\int_{\mathcal{S}\times\mathcal{A}}d_{h,\pi_{\mathrm{TS}}^{t}}^{\bar{\mathcal{E}}_{t}}(s,a)\cdot\frac{\Delta_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}(s,a)^{2}}{\alpha_{\mathcal{E}}^{2}}=\mathcal{I}^{t}.$ (A.12)
To summarize, by Eqn. (A.11) we have the following bound for $\Gamma_{t}^{\pi_{\mathrm{TS}}^{t}}$:
$\Gamma_{t}^{\pi_{\mathrm{TS}}^{t}}\leq\frac{\mathcal{I}^{t}\cdot\mathcal{T}^{t}}{\mathbb{I}_{t}^{\pi_{\mathrm{TS}}^{t}}\left(\widetilde{\mathcal{E}}_{t}^{*};(\mathcal{H}_{t},\mathcal{R}_{t,H})\right)}.$ (A.13)
@@
$\mathcal{I}^{t}$ $=\sum_{h=1}^{H}\mathbb{E}_{t}\mathbb{E}_{\pi_{\text{TS}}^{t}}^{\bar{\mathcal{E}}_{t}}\bigg{(}\mathbb{E}_{(s^{\prime},r^{\prime})\sim(P_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}\otimes R_{h}^{\widetilde{\mathcal{E}}_{t}^{*}})(\cdot|s_{h},a_{h})}\bigg{[}\frac{r^{\prime}+V_{h+1,\pi_{\mathcal{E}}^{*}}^{\widetilde{\mathcal{E}}_{t}^{*}}(s^{\prime})}{\alpha_{\mathcal{E}}}\bigg{]}-\mathbb{E}_{(s^{\prime},r^{\prime})\sim(P_{h}^{\bar{\mathcal{E}}_{t}}\otimes R_{h}^{\bar{\mathcal{E}}_{t}})(\cdot|s_{h},a_{h})}\bigg{[}\frac{r^{\prime}+V_{h+1,\pi_{\mathcal{E}}^{*}}^{\widetilde{\mathcal{E}}_{t}^{*}}(s^{\prime})}{\alpha_{\mathcal{E}}}\bigg{]}\bigg{)}^{2}$ (A.14)
$\overset{(a)}{=}\sum_{h=1}^{H}\mathbb{E}_{t}\mathbb{E}_{\pi_{\text{TS}}^{t}}^{\bar{\mathcal{E}}_{t}}\bigg{(}\mathbb{E}_{(s^{\prime},r^{\prime})\sim(P_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}\otimes R_{h}^{\widetilde{\mathcal{E}}_{t}^{*}})(\cdot|s_{h},a_{h})}\bigg{[}\frac{r^{\prime}-r_{h}^{\inf}(s_{h},a_{h})+V_{h+1,\pi_{\mathcal{E}}^{*}}^{\widetilde{\mathcal{E}}_{t}^{*}}(s^{\prime})-\inf_{s}V_{h+1,\pi_{\mathcal{E}}^{*}}^{\widetilde{\mathcal{E}}_{t}^{*}}(s^{\prime})}{\alpha_{\mathcal{E}}}\bigg{]}$
$\qquad-\mathbb{E}_{(s^{\prime},r^{\prime})\sim(P_{h}^{\bar{\mathcal{E}}_{t}}\otimes R_{h}^{\bar{\mathcal{E}}_{t}})(\cdot|s_{h},a_{h})}\bigg{[}\frac{r^{\prime}-r_{h}^{\inf}(s_{h},a_{h})+V_{h+1,\pi_{\mathcal{E}}^{*}}^{\widetilde{\mathcal{E}}_{t}^{*}}(s^{\prime})-\inf_{s}V_{h+1,\pi_{\mathcal{E}}^{*}}^{\widetilde{\mathcal{E}}_{t}^{*}}(s^{\prime})}{\alpha_{\mathcal{E}}}\bigg{]}\bigg{)}^{2}$
$\overset{(b)}{\leq}\frac{1}{2}\sum_{h=1}^{H}\mathbb{E}_{t}\mathbb{E}_{\pi_{\text{TS}}^{t}}^{\bar{\mathcal{E}}_{t}}\left[D_{\text{KL}}\left(\left(P_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}\otimes R_{h}^{\widetilde{\mathcal{E}}_{t}^{*}}\right)(\cdot|s_{h},a_{h})\middle\|\left(P_{h}^{\bar{\mathcal{E}}_{t}}\otimes R_{h}^{\bar{\mathcal{E}}_{t}}\right)(\cdot|s_{h},a_{h})\right)\right]$
$\overset{(c)}{\leq}\frac{1}{2}\mathbb{I}_{t}^{\pi_{\text{TS}}^{t}}\left(\widetilde{\mathcal{E}}_{t}^{*};(\mathcal{H}_{t},\mathcal{R}_{t,H})\right),$
where (a) adds and subtracts the constant term $r_{h}^{\inf}(s_{h},a_{h})$ and $\inf_{s}V_{h+1,\pi_{\mathcal{E}}^{*}}^{\widetilde{\mathcal{E}}_{t}^{*}}(s^{\prime})$, (b) uses the Pinsker’s inequality, (c) uses Lemma C.2. Plugging into Eqn. (A.13), we derive that $\Gamma_{t}^{\pi_{\text{TS}}^{t}}\leq\frac{1}{2}\mathcal{T}^{t}$, and thus
$BR_{T}(\pi_{\mathrm{IDS}})\leq\frac{1}{4\lambda}\mathbb{E}\left[\sum_{t=1}^{T}\mathcal{T}^{t}\right]+\frac{\lambda}{2}\log(K(\epsilon))+T\epsilon.$ (A.15)
Step 4 (Bound $\mathbb{E}[\mathcal{T}^{t}]$). The analysis tools used in this step is the Doob’s consistency theorem, with more details discussed in Appendix C.2. Define the true environment as $\mathcal{E}_{0}$. For brevity of notations, we define
@@
$\mathcal{T}^{t}=\sum_{h=1}^{H}\int_{(s,a)\in\mathcal{B}_{h,t}}\frac{\mathbb{E}_{t}\left[\alpha_{\mathcal{E}}^{2}\cdot d_{h,\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}(s,a)^{2}\right]}{\mathbb{E}_{t}\left[d_{h,\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}(s,a)\right]}\mathrm{d}\mu_{\mathcal{S}\times\mathcal{A}}.$ (A.16)
We now introduce another variable $\mathcal{E}^{\prime}$ to apply Lemma C.3. Let $\mathcal{E}^{\prime}\sim\mathbb{P}(\cdot|\mathcal{D}_{t})$ be a random variable independent of $\mathcal{E}$. By definition of $d_{h,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}}$, we have $\mathbb{E}_{t}\left[d_{h,\pi_{\mathcal{E}}^{*}}^{\bar{\mathcal{E}}_{t}}(s,a)\right]=\mathbb{E}_{\mathcal{E},\mathcal{E}^{\prime}\sim\mathbb{P}_{t}(\cdot)}\left[d_{h,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}^{\prime}}(s,a)\right]$, and
@@
$\leq\mathbb{E}_{\mathcal{E},\mathcal{E}^{\prime}\sim\mathbb{P}_{t}(\cdot)}\left[\alpha_{\mathcal{E}}^{2}\cdot d_{h,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}^{\prime}}(s,a)^{2}\right],$ (A.17)
@@
$\mathcal{T}^{t}\leq\sum_{h=1}^{H}\int_{\mathcal{S}\times\mathcal{A}}\frac{\mathbb{E}_{t}\left[\alpha_{\mathcal{E}}^{2}\cdot d_{h,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}^{\prime}}(s,a)^{2}\right]}{\mathbb{E}_{t}\left[d_{h,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}^{\prime}}(s,a)\right]}\chi_{\mathcal{B}_{h,t}}(s,a)\mathrm{d}\mu_{\mathcal{S}\times\mathcal{A}}=\sum_{h=1}^{H}\int_{\mathcal{S}\times\mathcal{A}}g_{t}(s,a,h,\mathcal{D}_{t})\mathrm{d}\mu_{\mathcal{S}\times\mathcal{A}},$ (A.18)
@@
This implies $g_{t}(s,a,h,\mathcal{D}_{t})\leq M_{d}H^{2}$ and $\mathcal{T}^{t}\leq M_{d}H^{3}$. By Lemma C.3, we have the following convergence results: for any $(s,a)\in\mathcal{S}\times\mathcal{A}$,
$\lim_{t\to\infty}\mathbb{E}_{t}\left[\alpha_{\mathcal{E}}^{2}d_{h,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}^{\prime}}(s,a)^{2}\right]=\alpha_{\mathcal{E}_{0}}^{2}d_{h,\pi_{\mathcal{E}_{0}}^{*}}^{\mathcal{E}_{0}}(s,a)^{2},$ (A.19)
$\lim_{t\to\infty}\mathbb{E}_{t}\left[d_{h,\pi_{\mathcal{E}}^{*}}^{\mathcal{E}^{\prime}}(s,a)\right]=d_{h,\pi_{\mathcal{E}_{0}}^{*}}^{\mathcal{E}_{0}}(s,a),$ (A.20)
@@
$\lim_{t\to\infty}\chi_{\mathcal{B}_{h,t}}(x)=\chi_{\mathcal{B}_{h}}(x),\qquad\text{where }\mathcal{B}_{h}\triangleq\left\{(s,a)\in\mathcal{S}\times\mathcal{A}\middle|\mathbb{E}\left[d_{h,\pi_{\mathcal{E}_{0}}^{*}}^{\mathcal{E}_{0}}(s,a)\right]\neq 0\right\}.$ (A.21)
From Eqn. (A.21), we can also derive that $\lim_{t\to\infty}\chi_{\mathcal{B}_{h,t}\bigcap\mathcal{B}_{h}}(x)=1$, and $\lim_{t\to\infty}\chi_{\mathcal{B}_{h,t}\setminus\mathcal{B}_{h}}(x)=0$. By Lebesgue dominated convergence theorem, we have